OpenStudy (anonymous):

You have 3 types of candies--sinckers, M&M's, and skittles In total you have 16 pieces of candy. The snickers weigh 8 grams each, the M&M's weigh 4 grams each, and the skittles 12 grams each. Together the bag weighs 124 grams you have 2 bags for skittles. How many of each do you have.

OpenStudy (anonymous):

i was able to come up with 2 equations....x+y+z+=16 and 8x+4y+12z=124 i just dont know how to make an equation out of "you have 2 bags of skittles"

OpenStudy (anonymous):

yea that's really random.. ignore that i guess

OpenStudy (anonymous):

ah that wouldn't work out well either..

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

ok ignore the last info...how will i do it when i ignore that?

OpenStudy (anonymous):

x + y + z = 16 2x + y + 3z = 31 (after simplified) x + 2z = 15 z = 15/2 - x/2 <--- x + y + z = 16 x + y + 15/2 - x/2 = 16 y = 17/2 - x/2 <--- Here you can get many solutions like x = 1, y = 8, z = 7 x = 3, y = 7, z = 6 x = 5, y = 6, z = 5 x = 7, y = 5, z = 4 x = 9, y = 4, z = 3 and so on

OpenStudy (anonymous):

but nothin is working :(

OpenStudy (anonymous):

we can assume it means there's 2 pieces of skittles lol x = 11, y = 3, z = 2 then

OpenStudy (anonymous):

haha :)...that's funny :P...the answer does look a little wierd because that means there is ONE piece of skittle in EACH bag

OpenStudy (anonymous):

the teacher wanted me to use substitution...everyone in my classes finished so quickly...i was just stuck

OpenStudy (anonymous):

pieces and bags <-- this confuses me too. i would just assume it means there's 2 pieces of skittles. that will make everything easier

OpenStudy (anonymous):

ok i will go with that :)...thank you so much for helping me :)